USE THE GRAPHING METHOD TO TELL HOW MANY SOLUTIONS THE SYSTEM HAS:
x+y=8
x+y=-1

Answer by ashleygurl92(6)

x+y=8
x+y=-1

multiply the bottom one by -1 to eliminate
x+y=8
-1(x+y=-1)
rewrite the problem
x+y=8
-x-y=-1
the x & y eliminates so tou're left with 0=9 which is NO SOLUTION because 0 doesn't equal 0... hope i helped!